Javier is currently paying ​$1,200 in interest on his credit cards annually.​ If, instead of paying​ interest, he saved this amount every​ year, how much would he accumulate in a​ tax-deferred account earning 8 percent over 11​, 16​, or 21 ​years?

If, instead of paying​ interest, this amount was saved every year in a​ tax-deferred account earning 8 percent over 11 ​years, the amount accumulated would be ​$?
Answer #1

We have to calculate the Future value of the sum Javier will have if he deposits a pmt of $1200 annually.

FV is calculated using the excel FV function:

11 years: =FV(8%,11,1200,,) = $19,974.5850

16 years: =FV(8%,16,1200,,) = $36,389.1397

21 years: =FV(8%,21,1200,,) = $60,507.5057